Given the vectors v(1) = <3,4>, v(2) = <1,1>, and v(3) = <2,7>
What is the magnitude of vector v(1)?

Answer by jim_thompson5910(35256)

Use the pythagorean theorem

a^2 + b^2 = c^2
3^2 + 4^2 = c^2
9 + 16 = c^2
c^2 = 25
c = sqrt(25)
c = 5

The magnitude of <3,4> is 5